12) Find the present value of an annuity that pays $3,000 per month for 16 years if money is worth 6%, compounded monthly.

can you please hand write the steps and the answer. please i need the answer.
Answer #1

Solution

Present value of annuity=Annuity amount*((1-(1/(1+r)^n))/r)

Where

r-periodic intrest rate-6/12=.5% per month

n=number of periods=16*12=192

Annuity amount=3000

Therefore

Present value of annuity=3000*((1-(1/(1+.005)^192))/.005)

Thus Present value of annuity=369714.0759